OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. R(3, 3), S(-6, -6) @some1

OpenStudy (anonymous):

Yay. okai so. do you know how to find slope with theeeseee?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

whatcha get for the slope? i'll check and make sure it's right...

OpenStudy (anonymous):

sureeeeee -.-

OpenStudy (anonymous):

YES GJ! it IS 1! :)

OpenStudy (anonymous):

is it..

OpenStudy (anonymous):

omg-.-

OpenStudy (anonymous):

lol yea. should i show you?

OpenStudy (anonymous):

fine..u had to disrupt me

OpenStudy (anonymous):

:O i'm sorry. tell meh. did you get how i did that?

OpenStudy (anonymous):

whell... u disrupted me.. what do i do next?

OpenStudy (anonymous):

-.- i'm sorry. did you get 1 using the slope formula?

OpenStudy (anonymous):

YES!

OpenStudy (anonymous):

3+6!!!!!!!!

OpenStudy (anonymous):

no. -9-3/-9-3...

OpenStudy (anonymous):

same thing..

OpenStudy (anonymous):

;D

OpenStudy (anonymous):

im soo mean..sorrry

OpenStudy (anonymous):

xD ok. anyway, So you would substitute that slope into ax+y=b. So right now, your equation should have 1x+y=b. Using one of the coordinate pairs given, substitute the values into the above equation and find b. Once you have the value of b, just put it back into the above equation! For example, 1(-6)+(-6)=b -12=b Sooooo: x+y=-12. der ya gooo :D and ish okai i don't mind if you're mean :) Comment if you don't get anythingggg

OpenStudy (anonymous):

:ooooooooooooooooooooooooooooooooh

OpenStudy (anonymous):

you don't get it? :(

OpenStudy (anonymous):

i got it.. hold on

OpenStudy (anonymous):

:Dthanks!!

OpenStudy (anonymous):

<3

OpenStudy (anonymous):

wait.. @some1.. so with these.. A(4, 1), B(5, 2) why is this answer wrong.. y=x-3

OpenStudy (anonymous):

okai so let's start from beginning. 2-1/5-4=1 so your slope is right. gj =) So it should look like this: 1x+y=b 1(4)+1=b 5=b x+y=5. der it isssss =)

OpenStudy (anonymous):

wait... so i cant take B(5, 2).. i have to do it with A?

OpenStudy (anonymous):

you can. look. 1(5)+2=b 7=b. x+y=7. You can do it with either. But problem is, your equation doesn't match either one :P

OpenStudy (anonymous):

oh..

OpenStudy (anonymous):

both answers are right?

OpenStudy (anonymous):

yush. i believe so.

OpenStudy (anonymous):

oh..

OpenStudy (anonymous):

lemme double check, i'm doubting myself too nao :P

OpenStudy (anonymous):

yep it works.

OpenStudy (anonymous):

ok:)

OpenStudy (anonymous):

Here use this and graph both. http://go.hrw.com/math/midma/gradecontent/manipulatives/GraphCalc/graphCalc.html The reason they are both on the same line is because of the slope. When the slope and the entire equation except b is exactly the same, the line is basically the same line.

OpenStudy (anonymous):

k :D

OpenStudy (anonymous):

lol. a medal for hard work :P

OpenStudy (anonymous):

hahaha ;D thanks